OpenStudy (anonymous):

Find the value of Sin(2sin-1(o.8))

OpenStudy (aravindg):

0.8 is same as 8/10=4/5 Hence sin-1 (o.8) will be sin-1 (4/5) =53 degree appprox

OpenStudy (aravindg):

2sin-1(4/5)=53*2

OpenStudy (aravindg):

So your question reduces to sin(53*2)=....

OpenStudy (anonymous):

How did you get 53*2 ?

OpenStudy (aravindg):

sin-1 (4/5) =53 2sin-1 (4/5) =2*53

OpenStudy (anonymous):

Am sorry but I dont understand. How did you get 53 in the first place ?

OpenStudy (aravindg):

Its elementary right triangle. The values are knowledge based.If you don't know that then use a calculator.

OpenStudy (anonymous):

Actually is there any other method ? we're not supposed to use calculators.

OpenStudy (aravindg):

okay use the elementary triangle

OpenStudy (aravindg):

 This is knowledge based data.You should keep this triangle in your mind.